OpenStudy (anonymous):

The y-intercept of the line whose equation is x - y = 7 is?

OpenStudy (zpupster):

put in form y=mx+b when x i s 0 b is the y intercept add y to both sides and subtract 7 both sides what do you get

OpenStudy (anonymous):

y=x+7

OpenStudy (zpupster):

x-y = 7 x-y+y = 7+y we added y both sides we get x=7+y now we subtract 7 x-7 = 7-7+y we get x-7=y or y = x-7 y-intercept is -7